Yost Rocks, Inc. just paid a $2 dividend, which it expects to maintain for the foreseeable future. The firm has a beta of 1.2 and U.S. Treasury bills yield 3%. If the market risk premium is 9.6%, what is the current price of the stock?


If the firm expects to maintain a constant 3% growth rate in dividends, what would the price be today?

If the firm was unable to pay a dividend for the next 4 years, but then paid a $2 dividend which grew at a constant 3%, what would the price be today? 


Nnswer: $13.77; $17.88; $10.09 10.
Answer #1

required return=risk free rate+beta*market risk premium=3%+1.2*9.6%=14.5200%

1.
P0=D0/(r-g)=2/(14.52%-3%)=17.3611111

2.
P0=D0*(1+g)/(r-g)=2*(1+3%)/(14.52%-3%)=17.8819444

3.
P0=D5/(1+r)^4*1/(r-g)=2/(1+14.52%)^4*1/(14.52%-3%)=10.0937411

> #1 is wrong

wiley Tue, Dec 7, 2021 8:38 PM